Question 2: Solve the following equation:

6 + log ⁡ ( 16 x 2 − 64 ) = log ⁡ ( 4 x + 8 ) + 8

A . x = 5.456

B . x = 3.859

C . x = 3.847

D . x = 27
E. none of the above
Choose the correct answer (A - E):
1
Expert's answer
2020-08-25T16:04:44-0400

6 + log ⁡ (16x2 − 64) = log ⁡ (4x + 8) + 8

Move terms

"log \u2061 (16x^2\u221264) - log (4x + 8) = 8 - 6"

"log \u2061 (16x^2 \u2212 64) - log (4x + 8) = 2"



From

"log (x) - log (y) = log (x\/y)"


"log \u2061 [(16x^2 \u2212 64)\/(4x + 8)] = 2"



Factor 4 out from expression.

"log \u2061 [4(4x^2 \u2212 16)\/4(x + 2)] = 2"


"log \u2061 [4x4(x^2 \u2212 4)\/4(x + 2)] = 2"


"log \u2061 [4(x^2 \u2212 4)\/(x + 2)] = 2"


Since from: "a^2 - b^2 = (a-b) (a+b)"


"log \u2061 [4(x \u2212 2)(x+2)\/4(x + 2)] = 2"


The expression simplifies to:

"log \u2061 [4(x-2)] = 2"


Since 2 = log 100

"log \u2061 [4(x-2)] = log100"

"[4(x-2)] = 100"

"4x-8= 100"

"4x= 108"

"x=27"



The answer is D = 27